OpenStudy (anonymous):

Assume all variables are implicit functions with respect to time. Find the indicated rate. y=(9-2x^2)^1/2 given y'=3 when x=2, find x'

OpenStudy (anonymous):

\[y'=\frac{-4xx'}{2\sqrt{9-2x^2}}\] is a start

OpenStudy (anonymous):

plug in the numbers you know to get the number you want

OpenStudy (anonymous):

\[3=-8x'/2\sqrt{1}\] ?

OpenStudy (anonymous):

did you really write \(\sqrt{1}\) ?

OpenStudy (anonymous):

i mean it is right, but not very helpful

OpenStudy (anonymous):

lol sorry its been a long day

OpenStudy (anonymous):

thanks i figured it out :)

OpenStudy (anonymous):

yw